# IrokoBench: A New Benchmark for African Languages in the Age of Large Language Models

### v1

@@ Abstract
Despite the widespread adoption of Large language models (LLMs), their remarkable capabilities remain limited to a few high-resource languages. Additionally, many low-resource languages (e.g., African languages) are often evaluated only on basic text classification tasks due to the lack of appropriate or comprehensive benchmarks outside of high-resource languages. In this paper, we introduce IrokoBench—a human-translated benchmark dataset for 16 typologically-diverse low-resource African languages covering three tasks: natural language inference (AfriXNLI), mathematical reasoning (AfriMGSM), and multi-choice knowledge-based QA (AfriMMLU). We use IrokoBench to evaluate zero-shot, few-shot, and translate-test settings (where test sets are translated into English) across 10 open and four proprietary LLMs. Our evaluation reveals a significant performance gap between high-resource languages (such as English and French) and low-resource African languages. We observe a significant performance gap between open and proprietary models, with the highest performing open model, Aya-101 only at 58% of the best-performing proprietary model GPT-4o performance. Machine translating the test set to English before evaluation helped to close the gap for larger models that are English-centric, like LLaMa 3 70B. These findings suggest that more efforts are needed to develop and adapt LLMs for African languages.
@@ 1 Introduction
In recent years, the capabilities of large language models (LLMs) have greatly improved, from coherent chat experiences to solving complex and knowledge-intensive tasks like mathematical reasoning, coding, and question answering [47, 33, 26]. These models have also demonstrated the ability to quickly learn new and challenging tasks with few in-context learning examples and through chain-of-thought reasoning [17, 53, 64]. However, most state-of-the-art LLMs are primarily trained on high-resource languages, resulting in sub-optimal performance for languages unseen during pre-training [58, 45]. Furthermore, this language coverage bias is reflected in the evaluation stage, predominantly conducted in English and a few other high-resource languages.
@@
There has been considerable effort to create benchmarks for African languages, but they typically cover simpler tasks, or are specific to narrow domains such as machine translation, and, more recently, reading comprehension [13, 11]. Some diverse reasoning benchmarks have included Swahili—the most spoken native African language—in benchmarks like commonsense reasoning [48] and natural language inference [21]. However, performance in Swahili does not reflect the performance on the over 2,000+ languages in Africa, which often lack substantial digital textual resources comparable to that available for Swahili. Consequently, current multilingual evaluations of LLMs do not accurately reflect capabilities in reasoning and knowledge-intensive tasks across the majority of African languages. Furthermore, the few comprehensive evaluations that exist across languages often rely on machine translation of English benchmarks [55]. While automatic translation is a popular approach given the cost and time investment required for human translation, it often suffers from noise and biases. For example, translated datasets often introduce artefacts from the translation model  [61, 28, 51] or fail to reflect cultural context appropriately [63, 32, 49]. Automatic curation may also amplify any of the ubiquitous issues with the quality of broad pretraining sets [39, 35, 24, 18].
@@
In this paper, we seek to address both the diversity and breadth of evaluation coverage. We introduce IrokoBench, a human curated benchmark dataset for 16 typologically diverse African languages which encompasses three complex tasks: natural language inference (NLI), mathematical reasoning, and multi-choice knowledge-based QA. The datasets were created by human translating a subset of English cross-lingual NLI (XNLI) [21], English Multilingual Grade School Math (MGSM) [54], and Massive Multitask Language Understanding (MMLU) [29], evaluation datasets into each of the 16 languages using professional translators.
@@
We conduct a large-scale evalution of IrokoBench to evaluate zero-shot, few-shot, and translate-test settings (where test sets are translated into English) across 10 open and four proprietary LLMs. Our main contributions can be enumerated as follows:
@@
    We introduce and release IrokoBench, a human-translated benchmark that includes languages from various geographical regions: six from West Africa, five from East Africa, four from Southern Africa, and one from Central Africa, all with varying degrees of “low-resourcedness.” [34].
@@
    Sharp cliff in performance across all models on low-resource languages Our evaluation shows a large gap ( $\sim$ $45$% on average) between the performance of high-resource languages (e.g., English) and African languages on all LLMs evaluated. Notably, Swahili performs better than other African languages, likely due to its frequent use as a benchmark.
@@
    Models generally perform poorly in in-language evaluation When evaluating models with prompts in native languages, state-of-the-art open models such as Gemma, Llama 3 are outperformed by massively multilingual models like Aya-101. This can be attributed to a lack of trained models to respond in the native languages of the users. Machine translating the test set to English before evaluation helped to close the gap for English-centric models; however, requiring users to always translate their prompts from their native language to English may not be desirable behavior.
@@
    IrokoBench highlights the performance divide between open and proprietary models on low-resource languages. We find that proprietary closed models generally outperform open models for African languages. However, even these proprietary models exhibit substantial performance drops, due to the limited monolingual web data for African languages. The lowest performance is observed in languages such as Ewe, Lingala, Luganda, Twi and Wolof, which each have less than 50 million characters of available data [36]. Among the tasks evaluated, AfriMGSM proves most challenging for LLMs, followed by AfriMMLU and AfriXNLI.
These results underline the need for focused development and adaptation of LLMs to better support African languages, especially those with limited data resources. We release IrokoBench on HuggingFace under the CC BY-SA 4.0 licence https://huggingface.co/collections/masakhane/irokobench-665a21b6d4714ed3f81af3b1 to further multilingual evaluation and research.
@@ 2 Related Work
The evaluation of multilingual capabilities of LLMs has garnered significant attention. This has led to an increase in research that explores their performance across diverse linguistic landscapes [7, 8, 37, 31, 14, 60, 55]. Despite this growing interest, there remains a notable lack of representation of African low-resource languages in these studies. [45, 12] address a broader spectrum of African languages, aligning more closely with our research. However, their study focuses on conventional NLP tasks, such as text classification, named entity recognition, question answering, and text generation. To address the lack of difficult benchmarks, a few works automatically translated MMLU benchmarks [47, 38, 60], but they do propagate errors of machine translation (MT) engines. Moreover, this is not applicable to low-resource languages with low-quality MT systems [2, 57]. Our research advances this by evaluating LLMs on more complex tasks using newly developed, human-annotated benchmarks specifically for African languages.
@@
Due to the limited representation of African languages in the field of NLP, there has been a growing effort to create benchmark datasets for African languages to enable research on these languages, especially evaluating existing NLP tools on these languages. Initiatives such as Masakhane have been instrumental in the creation of standard benchmark for tasks such as machine translation [2], named entity recognition [5, 3], part of speech tagging [23], news topic classification [6], and sentiment analysis [43]. There are also several multilingual benchmark datasets that cover a few African languages, such as SIB-200 [4], Flores [27, 57], Aya dataset and Collection [55] and Taxi1500 [40]. However, despite all these efforts, African languages still lack quality and more difficult datasets to evaluate the reasoning capabilities of LLMs, which have become ubiquitous. Hence, this work contributes to this space by creating datasets to evaluate general use LLMs in African languages in a variety of downstream tasks.
@@ 3 IrokoBench
No. of No. instances
Dataset languages Train/Dev/Test Subjects / Domains
AfriMGSM 15 (excl. Swahili) 8  /  -  /  250 grade school mathematics
AfriMMLU 17 (incl. French) 25  /  83  /  500 elementary mathematics, high-school geography, International law, global facts, high school microeconomics
AfriXNLI 15 (excl. Swahili) -  /  450  /  600 face-to-face, telephone, oxford university press (oup), fiction, travel, government, nineeleven, letters, slate, verbatim
3.1 Languages covered by IrokoBench
We cover 18 languages in our evaluation, 16 native African languages, and two European languages (English and French) which are widely spoken and official in many African countries. English and French are also the source languages we translated from. We cover 16 diverse and widely spoken African languages from four regions of Africa: six from West Africa (Ewe, Hausa, Igbo, Twi, Wolof, Yoruba), five from East Africa (Amharic, Kinyarwanda, Luganda, Swahili, and Oromo), four from Southern Africa (chiShona, isiXhosa, isiZulu, and Sesotho), and Central Africa (Lingala). The African languages are from two biggest language families in Africa: three from Afro-Asiatic and 13 from the Niger-Congo family—where we cover eight Bantu languages. Table 7 provides an overview of the languages covered, including their family, regions, and the number of native speakers.
3.2 Tasks covered by IrokoBench
The selection of these tasks is primarily driven by their coverage across various domains and downstream tasks for diverse use cases. Additionally, they enable the evaluation of logical, abstract, and reasoning capabilities in LLMs, which is the hallmark of human intelligence [16, 30]. Figure 1 provides examples of the different tasks covered in our datasets. We provide their descriptions below:
AfriXNLI
The task of NLI involves the classification of a pair of sentences—a premise and a hypothesis as to entailment, neutral, or contradiction semantic relation. For example, the sentence “so I’m not really sure why” contradicts “I am certain as to the reason why” but has a neutral relation to “I don’t know why he transferred schools”. Here, we human translate the English portion of XNLI (a multilingual dataset comprising 15 languages, including Swahili) into the 15 African languages (excluding Swahili). While the original XNLI dataset has over 2,500/5,000 as DEV/TEST split, Each language in AfriXNLI has only 450 DEV instances and 600 test instances. We selected an equal number of instances from the 10 domains of XNLI. The task is evaluated using the accuracy metric since the dataset has balanced classes.
@@ 3.2 Tasks covered by IrokoBench
This is a multi-choice knowledge QA curated from freely available online sources by undergraduate and graduate students in the USA. The subjects cover simple general knowledge questions like “global fact” to highly-technical questions like “professional law” and “professional medicine”. MMLU are often grouped into STEM, Humanities, Social Sciences, and Others category. We focus on five subjects that we believe present less tendency to be culturally unbiased and that are not too difficult to translate since many of the subjects covered are only taught in African countries using English or French, making it extremely difficult to translate highly technical subjects, especially the STEM subjects. Table 1 shows the five subjects covered: two social science subjects (high-school geography and high-school microeconomics), one STEM subject (elementary mathematics), one humanities subject (international law), and one OTHER category (global facts). In total, we translated 608 question-answer pairs, with 500 instances in the test split, 100 questions per subject. The task is evaluated using the option prediction accuracy.
@@
This is a QA task with questions obtained from grade school mathematical word problems created by human problem writers. AfriMGSM expands the original MGSM dataset [54], which contains 250 QA pairs and 11 languages (including Swahili), to 15 more languages. The dataset consists of 8 training examples for few-shot and chain-of-thought prompting and 250 as a test set. We evaluated this task using the Exact Match metric, which is popularly used for QA tasks.
@@ 3.3 Data collection process
We recruited language coordinators for each of the 16 African languages and French, and asked them to recruit professional translators to translate the sentences. The translation process took about two months, they started with XNLI, then MGSM and MMLU. Each translator received an appropriate remuneration for their work.  [^1] We recruited a logistic company in Kenya that managed all recruitment and payments—each country has different rates. Most of the translators translated from English except for Ewe, Lingala and Wolof translators that translated from French since they are from the Francophone region of Africa. Additionally, we translated the MMLU dataset to French by professional translators and from French to these three languages. Many of the Francophone translators understand French and English but are more fluent in French, so they could cross-check from English if the French sentences were not clear enough.
@@
Regarding quality control, language coordinators reviewed and corrected any poorly translated sentences. Translators received payment only after this phase to ensure the quality of translations. For additional checks, we computed COMET [50] quality estimation (QE) scores between the human translation and the original sentences based on AfriCOMET QE metric [62]. In general, the distribution of the scores (range between 0 and 1) reflect that most translated sentences are between $0.7$ and $1.0$ for about 13 language pairs except for Lingala, Twi, and Wolof where the average is around $0.5$. Further analysis shows that we cannot rely on these scores for those three languages since they are not covered in the pre-training of the original AfroXLMR encoder [9] used to build the AfriCOMET QE metric. Similar findings were reported in the original AfriCOMET QE paper that Twi had worse correlation with human judgement (i.e., 0.279 for Pearson, and 0.060 for Spearman) [62]. We provide further analysis of the COMET scores in Appendix A.3.

### v2

@@ Abstract
Despite the widespread adoption of Large language models (LLMs), their remarkable capabilities remain limited to a few high-resource languages. Additionally, many low-resource languages (e.g., African languages) are often evaluated only on basic text classification tasks due to the lack of appropriate or comprehensive benchmarks outside of high-resource languages. In this paper, we introduce IrokoBench—a human-translated benchmark dataset for 17 typologically-diverse low-resource African languages covering three tasks: natural language inference (AfriXNLI), mathematical reasoning (AfriMGSM), and multi-choice knowledge-based question answering (AfriMMLU). We use IrokoBench to evaluate zero-shot, few-shot, and translate-test settings (where test sets are translated into English) across 10 open and six proprietary LLMs. Our evaluation reveals a significant performance gap between high-resource languages (such as English and French) and low-resource African languages. We observe a significant performance gap between open and proprietary models, with the highest performing open model, Gemma 2 27B only at 63% of the best-performing proprietary model GPT-4o performance. In addition, machine translating the test set to English before evaluation helped to close the gap for larger models that are English-centric, such as Gemma 2 27B and LLaMa 3.1 70B. These findings suggest that more efforts are needed to develop and adapt LLMs for African languages.
IrokoBench: A New Benchmark for African Languages in the Age of Large Language Models
David Ifeoluwa Adelani1,2∗, Jessica Ojo1,3∗, Israel Abebe Azime4∗, Jian Yun Zhuang5, Jesujoba O. Alabi4, Xuanli He6, Millicent Ochieng7, Sara Hooker8, Andiswa Bukula9, En-Shiun Annie Lee10, Chiamaka Chukwuneke11, Happy Buzaaba12, Blessing Sibanda, Godson Kalipe, Jonathan Mukiibi13∗, Salomon Kabongo14∗, Foutse Yuehgoh15∗, Mmasibidi Setaka9, Lolwethu Ndolela, Nkiruka Odu, Rooweither Mabuya9, Shamsuddeen Hassan Muhammad16, Salomey Osei17∗, Sokhar Samb18∗, Tadesse Kebede Guge19∗, Tombekai Vangoni Sherman20, Pontus Stenetorp6 Masakhane NLP, 1Mila, McGill University, 2Canada CIFAR AI Chair, 3Lelapa AI, 4Saarland University, 5University of Toronto, 6University College London, 7Microsoft Research Africa, 8Cohere For AI, 9SADiLaR, 10Ontario Tech University, 11Lancaster University, 12Princeton university, 13Makerere University, 14Leibniz Universität Hannover, 15Le CNAM, 16Imperial College London, 17Universidad de Deusto, 18DAUST, 19Haramaya University.
@@ 1 Introduction
In recent years, the capabilities of large language models (LLMs) have greatly improved, from coherent chat experiences to solving complex and knowledge-intensive tasks like mathematical reasoning, coding, and question answering (QA) (OpenAI et al., 2024; Jiang et al., 2024; Gemini-Team et al., 2024). These models have also demonstrated the ability to quickly learn new and challenging tasks with few in-context learning examples and through chain-of-thought reasoning (Brown et al., 2020; Shi et al., 2022; Wei et al., 2022). However, most state-of-the-art (SoTA) LLMs are primarily trained on high-resource languages (HRLs), resulting in sub-optimal performance for languages unseen during pre-training Touvron et al. (2023); Ojo et al. (2023). Furthermore, this language coverage bias is reflected in the evaluation stage, predominantly conducted in English and a few other HRLs.
@@
There has been considerable effort to create benchmarks for African languages, but they typically cover simpler tasks, or are specific to narrow tasks such as machine translation, and, more recently, reading comprehension (Bandarkar et al., 2023; Aremu et al., 2023). Some diverse reasoning benchmarks have included Swahili—the most spoken native African language—for tasks like commonsense reasoning (Ponti et al., 2020) and natural language inference (Conneau et al., 2018). Consequently, current multilingual evaluations of LLMs do not accurately reflect their capabilities in reasoning and knowledge-intensive tasks across the majority of African languages.
Furthermore, the few comprehensive evaluations that exist across languages often rely on machine translation of English benchmarks (Singh et al., 2024). While automatic translation from English benchmarks is a popular approach given the cost and time investment required for human translation, it often suffers from noise and biases  (Vanmassenhove et al., 2021; Lee et al., 2022; Khiu et al., 2024; Hartung et al., 2023; Savoldi et al., 2021) or fail to reflect cultural context (Wang et al., 2022; Ji et al., 2023; Pudjiati et al., 2022). Automatic curation may also amplify any of the ubiquitous issues with the quality of broad pretraining sets (Luccioni and Viviano, 2021; Kreutzer et al., 2022; Ferrara, 2023).
@@
In this paper, we seek to address both the diversity and breadth of evaluation coverage. We introduce IrokoBench, a human curated benchmark dataset for 17 typologically diverse African languages which encompasses three complex tasks: natural language inference (NLI), mathematical reasoning, and multi-choice knowledge-based QA. The datasets were created by human translating a subset of English cross-lingual NLI (XNLI) (Conneau et al., 2018), English Multilingual Grade School Math (MGSM) (Shi et al., 2023), and Massive Multitask Language Understanding (MMLU) (Hendrycks et al., 2021a), evaluation datasets into each of the 16 languages using professional translators.
@@
We conduct a large-scale evaluation of IrokoBench to assess zero-shot, few-shot, and translate-test settings (where test sets are translated into English) performance across 10 open and six proprietary LLMs. Our main contributions can be enumerated as follows:
@@
    We introduce and release IrokoBench, a human-translated benchmark that includes 16 languages from various geographical regions in Africa, all with varying degrees of “low-resourcedness” (Joshi et al., 2020).
@@
    Sharp cliff in performance across all models on low-resource languages Our evaluation shows a large gap ( $\sim$ $45$% on average) between the performance of high-resource languages (e.g., English) and African languages on all LLMs evaluated. Notably, Swahili performs better than other African languages, likely due to its large corpus on the web.
@@
    Models generally perform poorly in in-language evaluation This can be attributed to the inability of current SoTA LLMs to respond in the native languages of the users. Machine translating the test set to English before evaluation helped to close the gap for English-centric models; however, requiring users to always translate their prompts to English may not be a desirable behavior.
@@
    IrokoBench highlights the performance divide between open and proprietary models on low-resource languages. We find that proprietary closed models generally outperform open models for African languages. However, even these proprietary models exhibit substantial performance drops, due to the limited monolingual web data for African languages. The lowest performance is observed in languages such as Ewe, Lingala, Luganda, Twi and Wolof, which each have less than 50 million characters of available data (Kudugunta et al., 2023). Among the tasks evaluated, AfriMGSM proves most challenging for LLMs, followed by AfriMMLU and AfriXNLI.
We will release IrokoBench on GitHub under the CC BY-SA 4.0 licence upon acceptance.
@@ 2 Related Work
The evaluation of multilingual capabilities of LLMs has garnered significant attention. This has led to an increase in research that explores their performance across diverse linguistic landscapes Ahuja et al. (2023a, b); Lai et al. (2023b); Hendy et al. (2023); Bang et al. (2023); Üstün et al. (2024); Singh et al. (2024). Despite this growing interest, there remains a notable lack of representation of African low-resource languages in these studies. Ojo et al. (2023); Azime et al. (2024) address a broader spectrum of African languages, aligning more closely with our research. However, their study focuses on conventional NLP tasks, such as text classification, question answering, and text generation tasks. To address the lack of difficult benchmarks, a few works automatically translated MMLU benchmarks (Lai et al., 2023a), but they do propagate errors of machine translation (MT) engines. Moreover, this is not applicable to low-resource languages with low-quality MT systems Adelani et al. (2022a); Costa-jussà et al. (2024). Our research advances this by evaluating LLMs on more complex tasks using newly developed, human-annotated benchmarks specifically for African languages.
@@
Due to the limited representation of African languages in the field of NLP, there has been a growing effort to create benchmark datasets for African languages to enable research on these languages. Initiatives such as Masakhane have been instrumental in the creation of standard benchmark for tasks such as machine translation Adelani et al. (2022a), named entity recognition Adelani et al. (2021, 2022b), part of speech tagging Dione et al. (2023), news topic classification Adelani et al. (2023), and sentiment analysis Muhammad et al. (2023). There are also several multilingual benchmark datasets that cover a few African languages, such as SIB-200 Adelani et al. (2024), Flores Goyal et al. (2022); Costa-jussà et al. (2024), Aya dataset and Collection (Singh et al., 2024) and Taxi1500 Ma et al. (2023). However, despite all these efforts, African languages still lack quality and more difficult datasets, our work fills this gap.
@@ 3 IrokoBench
3.1 Languages covered by IrokoBench
IrokoBench  [^1] Our benchmark name comes from Ìrókò— a large hardwood in West Africa that is very durable for making bench. cover 19 languages, 17 native African languages, and two European languages (English and French) which are widely spoken and official in many African countries. English and French are also the source languages we translated from. The 17 diverse and widely spoken African languages are from four regions of Africa: seven from West Africa (Ewe, Hausa, Igbo, Twi, Vai, Wolof, Yoruba), five from East Africa (Amharic, Kinyarwanda, Luganda, Swahili, and Oromo), four from Southern Africa (chiShona, isiXhosa, isiZulu, and Sesotho), and Central Africa (Lingala). These languages are from three language families in Africa: one from Mande, three from Afro-Asiatic and 13 from the Niger-Congo family—where we cover eight Bantu languages. Table 9 provides an overview of the languages covered, including their family, regions, and the number of native speakers.
3.2 Tasks covered by IrokoBench
The selection of these tasks is primarily driven by their coverage across various domains and downstream tasks for diverse use cases. Additionally, they enable the evaluation of logical, abstract, and reasoning capabilities in LLMs, which is the hallmark of human intelligence Bowman et al. (2015); Hendrycks et al. (2021b). Figure 1 provides examples of the different tasks covered in our datasets. We provide their descriptions below:
AfriXNLI
The task of NLI involves the classification of a pair of sentences—a premise and a hypothesis as entailment, neutral, or contradiction semantic relation. For example, the sentence “so I’m not really sure why” contradicts “I am certain as to the reason why” but has a neutral relation to “I don’t know why he transferred schools”. Here, we human translate the English portion of XNLI (a multilingual dataset comprising 15 languages, including Swahili) into the 15 African languages (excluding Swahili). While the original XNLI dataset has over 2,500/5,000 as DEV/TEST split, Each language in AfriXNLI has only 450 DEV instances and 600 test instances. We selected an equal number of instances from the 10 domains of XNLI. The task is evaluated using the accuracy metric since the dataset has balanced classes.
No. of No. instances Average Length
Dataset languages Train/Dev/Test Subjects / Domains Train/Dev/Test
AfriMGSM 17 (excl. Swahili, inc. Vai) 8  /  -  /  250 grade school mathematics 25  /  -  /  46
AfriMMLU 17 (incl. French) 25  /  83  /  500 elementary mathematics, high-school geography, International law, global facts, high school microeconomics 18  /  17  /  17
AfriXNLI 15 (excl. Swahili) -  /  450  /  600 face-to-face, telephone, oxford university press (oup), fiction, travel, government, nineeleven, letters, slate, verbatim -  / 10  / 10 (hyp.)   -  / 18  / 17 (pre.)
@@ 3.2 Tasks covered by IrokoBench
This is a multi-choice knowledge QA curated from freely available online sources by undergraduate and graduate students in the USA. The subjects cover simple general knowledge questions like “global fact” to highly-technical questions like “professional law” and “professional medicine”. MMLU are often grouped into STEM, Humanities, Social Sciences, and Others category. We focus on five subjects that we believe are culturally unbiased (or international) and that are simpler to translate since many of the subjects covered are only taught in African countries using English or French, making it extremely difficult to translate highly technical subjects, especially the STEM subjects. Table 1 shows the five subjects covered: two social science subjects (high-school geography and high-school microeconomics), one STEM subject (elementary mathematics), one humanities subject (international law), and one OTHER category (global facts). In total, we translated 608 question-answer pairs, with 500 instances in the test split, 100 questions per subject. The task is evaluated using the option prediction accuracy.
@@
This is a QA task with questions obtained from grade school mathematical word problems created by human problem writers. AfriMGSM expands the original MGSM dataset (Shi et al., 2023), which contains 250 QA pairs and 11 languages (including Swahili), to 15 more languages. The dataset consists of 8 training examples for few-shot and chain-of-thought prompting and 250 as a test set. We evaluated this task using the Exact Match metric, which is popularly used for QA tasks.
@@ 3.3 Data collection process
We recruited language coordinators for each of the 16 African languages and French, and asked them to recruit professional translators to translate the sentences. The translation process took about two months, they started with XNLI, then MGSM and MMLU. Each translator received an appropriate remuneration for their work.  [^2] We recruited a logistic company in Kenya that managed all recruitment and payments—each country has different rates. For example, we paid \$549.78 for the translation of 1020 XNLI samples in South Africa, \$355.86 in Nigeria. Most of the translators translated from English except for Ewe, Lingala and Wolof translators that translated from French since they are from the Francophone region of Africa. Additionally, we translated the MMLU dataset to French by professional translators and from French to these three languages. Many of the Francophone translators understand French and English but are more fluent in French, so they could cross-check from English if the French sentences were not clear enough.
@@
Regarding quality control, language coordinators reviewed and corrected any poorly translated sentences. Translators received payment only after this phase to ensure the quality of translations. For additional checks, we computed COMET (Rei et al., 2020) quality estimation (QE) scores between the human translation and the original sentences based on AfriCOMET QE metric (Wang et al., 2024). In general, the distribution of the scores (range between 0 and 1) reflect that most translated sentences are between $0.7$ and $1.0$ for about 13 language pairs except for Lingala, Twi, and Wolof where the average is around $0.5$. Further analysis shows that we cannot rely on these scores for those three languages since they are not covered in the pre-training of the original AfroXLMR encoder (Alabi et al., 2022) used to build the AfriCOMET QE metric. Similar findings were reported in the original AfriCOMET QE paper that Twi had worse correlation with human judgement (i.e., 0.279 for Pearson, and 0.060 for Spearman) (Wang et al., 2024). We provide further analysis of the COMET scores in Appendix A.3.